ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

MSC2000

1991 MS Classification: 20 Kister, Jane, Mathematical Reviews, USA; Wegner, Bernd, Zentralblatt für Mathematik, Germany The year 2000 Mathematics Subject Classification, MSC2000 Since 1996, the editors of Mathematical Reviews (MR) and Zentralblatt für Mathematik (ZBL) have been collaborating on a revision of the 1991 Mathematics Subject Classification (MSC). It is planned that the revised classification, MSC2000, will be used by the two journals starting in the year 2000. We will present the new classification, in particular describing the principles underlying the revision and the major changes in the MSC.